OpenStudy (anonymous):

Can someone tell me what commutaive property, associative property and identiy prep means??

OpenStudy (anonymous):

In mathematics an operation is \(\Large\underline{commutative}\) if changing the order of the operation does not change the end result. Think of this property as it sounds "commute" A couple of basic examples would be that a x (b x c) = (a x b) x c or... a + (b+ c)=(a + b) + c though the factors are grouped differently, they still come out the same. In mathematics, the \(\Large\underline{associative\ property}\) is a property of some binary operations. In propositional logic, associativity is a valid rule of replacement for expressions in logical proofs. Within an expression containing two or more occurrences in a row of the same associative operator, the order in which the operations are performed does not matter as long as the sequence of the operands is not changed. That is, rearranging the parentheses in such an expression will not change its value. Consider, for instance, the following equations: (5+2)+1=5+(2+1)=8 5 x (5 x 3)= (5 x 5) x 3=75 \(\Large\underline{Identity\ property}\) of addition states that the sum of zero and any number or variable is the number or variable itself. For example, 4 + 0 = 4, - 11 + 0 = - 11, y + 0 = y are few examples illustrating the identity property of addition. Identity property of multiplication states that the product of 1 and any number or variable is the number or variable itself. For example, 4 × 1 = 4, - 11 × 1 = - 11, y × 1 = y are few examples illustrating the identity property of multiplication. \(\Huge\bf\color{darkorchid}{I\ hope\ that\ helped!\ :D}\)